# NeuroPilot: A Realtime Brain-Computer Interface system to enhance concentration of students in online learning

### v1

@@ Abstract
Prevalence of online learning poses a vital challenge in real-time monitoring of students’ concentration. Traditional methods such as questionnaire assessments require manual interventions and webcam-based monitoring fails to provide accurate insights into learners’ mental focus as they are deceived by mere screen fixation without cognitive engagement. Existing BCI-based approaches lack real-time validation and evaluation procedures. To address these limitations, a Brain-Computer Interface (BCI) system is developed using a non-invasive Electroencephalogram (EEG) headband, FocusCalm, to record brainwave activity under attentive and non-attentive states. 20 minutes of data were collected from each of 20 participants watching a pre-recorded educational video. The data validation employed a novel intra-video questionnaire assessment. Subsequently, collected signals were segmented (sliding window), filtered (butterworth bandpass), and cleaned (removal of high-amplitude and EOG artifacts such as eye blinks). Time, frequency, wavelet and statistical features have been extracted, followed by recursive feature elimination (RFE) with Support vector machines (SVMs) to classify attention and non-attention states. The leave-one-subject-out (LOSO) cross-validation accuracy has been tested to be 88.77%. The system provides feedback alerts upon non-attention state detection and keeps focus profile logs. A pilot study was conducted to evaluate the effectiveness of real-time feedback. Five participants completed a 10-minute session consisting of a 5-minute baseline phase without feedback followed by a 5-minute feedback phase, during which alerts were issued if participants remained non-attentive for approximately 8 consecutive seconds. A paired t-test (t = 5.73, p = 0.007) indicated a statistically significant improvement in concentration during the feedback phase.
@@ I Introduction
Concentration monitoring has become essential to ensure learning effectiveness in online classes, whose adoption has been steadily increasing across different educational curricula [1]. Prior research shows that learners are more prone to mind-wandering when they are watching lecture videos (being regarded as online learning scenario) than during live lectures (traditional offline classes) [2]. Thus, to track the engagement of students in educational and learning settings, conventional methods of developing questionnaires [3] or camera-based monitoring [4, 5] are used. Although questionnaires are useful for objective evaluation of students, they only serve as a manual tool that requires human intervention and input. Conversely, webcam-based monitoring systems are limited to analyzing external behavioral cues such as facial expressions, gaze, or posture, and therefore lack the ability to provide direct neurophysiological insights into learners’ cognitive attention levels [6], alongside inducing privacy issues [7]. Additionally, vision-based models fail to preserve robustness in low-light conditions. For this reason, Brain-Computer Interface (BCI) has emerged as a powerful tool to perform cognitive load evaluation of students while watching lecture videos.
@@
BCI is a form of neurotechnology that is leveraged to establish a direct connection between humans and machines. Specifically, electroencephalography (EEG) signals are increasingly being utilized for various use cases, including bionic intelligence[8], vehicle control [9, 10, 11], and other assistive technologies [12, 13, 14]. BCI has gained considerable recognition in the classification of attentive and non-attentive states [15] including distraction [16] and drowsiness[17]. Despite the fact that this article is particularly oriented to enhance concentration levels, terms like focus, attention, concentration, and engagement, which are different constructs overlapping each other, can be used interchangeably as the goal is to improve online learning for students using BCI. It has been seen from studies that the prefrontal cortex plays a crucial role in evaluating sustained attention [18, 19, 20]. In the frontal region, the features contributing to concentration level detection are predominantly the changes in delta, theta, alpha, and beta band power. [21, 18, 19].
@@
Current state-of-the-art works use EEG devices for laboratory experimentation that are not fully consumer-grade which makes them not suitable for real-time implementation. Some of them are Emotiv Insight, Emotiv EPOC+, and OpenBCI Ultracortex. These devices lack dry electrodes, requires accurate placement across multiple scalp regions, not cost-effective and present setup challenges for consumers [22, 23]. Among consumer-grade devices, cost-effective options such as Muse, NeuroSky MindWave, and FocusCalm have been studied. However, existing works provide limited details about their data collection strategies in online learning scenarios [24, 25, 26, 27], leaving a gap in collecting ground truth (referring to correct labeling of neural data containing actual concentration and actual non-concentration states). Other consumer-grade devices from companies such as Cognionics and Wearable Sensing have been promoted for use in sports scenarios but their high cost and large number of electrodes reduce their suitability for consumer-friendly applications in this field. Vital gaps are seen in real-time implementation of BCI research articulated to track concentration of students in learning environments. Huang et al. [28] have focused on real-time cognitive monitoring to improve concentration time with neuro-feedback mechanism. However, the data collection strategy included mental arithmetic tasks which is not an identical simulation of learning environment and the data validation has been done using post regulation assessments such as Sustained Attention to Response Task (SART) [29], which cannot validate the participants’ engagement during the mental arithmetic tasks. Furthermore, the working algorithm of the single-channel EEG headband, along with the feature extraction and model training process, has not been discussed, as their work focuses more on evaluating the effects of neurofeedback rather than on developing the classification methodology.Conrad and Newman have developed a strategy to use EEG signals for evaluating mind-wandering in online classes [30]. However, the use of only pre- and post-assessment questionnaires is insufficient to validate learners’ attention levels during the video. The work done by Rehman et al. [31] relied on publicly available datasets which poses the same limitations as the earlier research gaps. The studies summarized in Table I were selected to represent existing approaches for monitoring learner attention in online learning environments. Inclusion criteria were works that addressed learner engagement or attention in online or remote learning settings, and use of other related modalities (e.g., webcam-based, neurofeedback) whose limitations are being addressed in our proposed approach. Exclusion criteria included works focusing purely on offline classroom learning, cognitive tasks unrelated to learning (e.g., driving simulators), or studies that did not explicitly address attention/engagement measurement.
@@
In this study, a eeg based attention classification framework has been proposed to address the research gaps, with a focus on real-time implementation in online learning scenarios. A novel data acquisition and validation design through intra-video questionnaire assessment has been done to label the blocks as ground truths. The experimentation has been done on classifying two cognitive states of students, namely attention and non-attention. Twenty participants were recruited to participate in data collection experiments containing two sessions, designated for the two classes of data and the device used is a FocusCalm headband. After that, several preprocessing techniques have been employed to get clean data, followed by feature extraction, feature selection and classification. Following this, leave-one-subject-out validation has been done in order to evaluate the classifier model when faced with unseen data. Finally a simple app GUI has been developed to have real-time insights of concentration levels along with a feedback mechanism to alert the learner during non-attention states. Thus, this article has contributions in the following areas:
  • A data collection paradigm has been introduced that uses authentic lecture videos from educational technology platforms as EEG stimuli, offering a closer approximation to real-world online learning environments than traditional controlled tasks.
  • A novel intra-video questionnaire assessment has been proposed for data validation, which, unlike traditional post-task questionnaires or external measures (e.g., SART), collects learner responses during the lecture itself. This approach minimizes recall bias, provides time-specific ground-truth labels, and enhances the reliability of the EEG dataset.
  • A real-time system has been developed with feedback alerts to ensure the attention spans of learners improve over time. A preliminary evidence of its effectiveness in enhancing concentration is validated via a pilot study.
@@ II Data Collection
In this section, a detailed explanation of the inclusion criteria of participants, device used for data collection, experimental design and setup is given.
@@ II-A Participant Selection
A total of 20 students with a mean age of $22.8\pm 3.8$ years have been recruited for the data collection experiment. A detailed explanation of the experimental setup, collection paradigm and the study has been given to the participants before taking their informed consents. Participants retained the right to leave the data collection process at their discretion at any moment. This data collection protocol has been approved by the Institutional Research Ethics Board (IREB), United International University (UIU). Among the participants, 15 are graduate students of Advanced Intelligent Multidisciplinary System Lab (AIMS Lab) under Institute of Research, Innovation, Incubation & Commercialization (IRIIC), UIU and rest are the students of UIU. They have been selected to pass the condition of having cognitive abilities necessary to comprehend standard high school or equivalent curriculum-based educational material. This has been ensured due to the video stimuli being exposed to the participants at the time of data collection containing online lectures on different topics prepared by national educational technology platforms.
@@ II-B Data Acquisition Device
FocusCalm headband [32] has been used as the data acquisition device in this work for its consumer-grade accessibility, dry electrode system, frontal EEG coverage, low setup complexity and a balanced trade-off between signal quality and user comfort. It is a neuro-feedback EEG device designed for meditation purposes. There are three dry electrodes placed close to each other and aligns with the Fp1, Fp2 and Fpz positions of the international 10-10 electrode placement system as shown in Fig. 1. Fpz is used as the primary electrode, while Fp1 and Fp2 serve as the reference and ground electrodes [33]. The device wirelessly sends single channel data at 250 Hz along with eight distinct values of which six indicate the delta, theta, alpha, low beta, high beta and gamma frequency band values and the other two shows attention and meditation scores given by their developed model. The devlopers of FocusCalm at BrainCoTech has made a software development kit (SDK) named Crimson SDK which is accessible for taking raw eeg data along with the eight other values in real-time.
@@ II-C Data Collection Experiment
The data collection experiment has been designed using the Psychopy software [34]. The design included a starting instructional video for both classes (attentive and non-attentive) that gives a brief view of the experiment done.
The participants are asked to watch a 10 minutes long video lecture that has been splitted into blocks of 30 seconds mean duration. Subsequently, a question on that exact block will be showed on the screen with multiple choices which the participants have to answer. The duration given for answering the question has been set to 15 seconds as participants take around 4 to 5 seconds for reading a sentence [35]. This gives the participants 10 seconds to answer which is ample considered that humans take around 2 seconds to recognize answers [36]. The questions have been set to ensure that the participant has followed the lecturer’s quotes and drawings on the board. The questions require no external knowledge to be answered in order to ensure objective evaluation of a participant’s engagement. After the instructional video, a short sample video is played followed by a question answer block to make the participant get accustomed to the experiment. A conditional screen is played with a fixation cross where the participant has to press a key to start the video lecture. In this way, each participant had to watch 20 video blocks of a lecture followed by questions inserted between them. The experimental setup captures timestamps which mark the start and end of each video block together with participants’ responses to the questions. These records are subsequently used for data validation.
For the non-attentive data collection session, similar video sequences are presented but rendered inaudible and blurred to prevent participants from perceiving the content. The participants were instructed to stare at the screen for 5 minutes followed by 5 minutes with eyes closed. In the eyes-open condition, participants viewed the hazy, non-informative video intended to provide minimal external stimulation and thereby facilitate mind-wandering, a common occurance in real-world scenarios where students face difficulties in maintaining attention during online lectures. The eyes-closed condition has been designed to simulate drowsiness or dozing off during online lectures. A brief overview has been illustrated in Fig. 2a.
@@ II-D Data Validation
After succesful data acquistion, the eeg data are stored with proper timestamps through some modifications of the Crimson SDK and the responses to the questions are stored with synchronized timestamps. The correct response to a question from the participants validates the EEG data corresponding to the video block that the question succeeds. All the validated blocks are stored as ground truths for attentive data. For non-attentive data, the questionnaire assessment is not done and thus require no further validation. The procedure has been shown in Fig. 2b.
@@ III Methods
The validated data is processed through several steps namely data pre-processing, feature extraction and feature selection before being used as training data for classification.
@@ III-A Data Preprocessing and Preparation
Data is pre-processed through steps that include segmentation, filtering, noise removal and smoothing.
@@ III-A1 Segmentation
EEG data of each participant of a total length of $T$ samples is partitioned to get data segments of optimal length $L=1750$ samples (details in section III.D) . Although the average sampling frequency of 250 Hz is retained, samples have been considered instead of time intervals due to minute temporal delays in packet transmission that complicate timestamp-based calculations. Through validation, $K$ number of blocks are selected from the initially acquired $n$ blocks for each participant. This number corresponds to the average number of questions (around 20) set for a video lecture per session. Sliding window has been used during segmentation to reduce the edge effects and improve statistical reliability [37]. The sliding window is also referred in this article through overlap ratio $r=0.7$ (details in section III.D). The total number of segments found for each selected block is:
@@
And the total number of samples are $x=n\times L$. During segmentation, two columns have been added for labeling and assigning subject-wise user_ids. These will be used later in classification.
To ensure subject-wise label balance, samples with both labels have been extracted separately for each subject. Let $\mathcal{D}_{u}^{0}$ and $\mathcal{D}_{u}^{1}$ denote the sets of samples with labels 0 and 1 respectively (0 meaning attentive and 1 meaning non-attentive), for a given user $u$. An equal number of samples are retained by trimming the larger class to match the size of the smaller one. Number of samples to be retained, $n_{u}=\min(|\mathcal{D}_{u}^{0}|,|\mathcal{D}_{u}^{1}|)$, and the last $n_{u}$ samples from each label set are selected, which gives $\mathcal{D}_{u,\text{tail}}^{0}(n_{u})$ and $\mathcal{D}_{u,\text{tail}}^{1}(n_{u})$. The balanced dataset for user $u$ is then
@@
Given that the dataset retains a substantial number of segments, class imbalance is addressed by trimming the larger class rather than oversampling the smaller one.

### v2

# NeuroPilot: A Real-time Brain-Computer Interface system to enhance concentration of students in online learning

@@ Abstract
The prevalence of online learning poses a vital challenge in real-time monitoring of students’ concentration. Traditional methods such as questionnaire assessments require manual intervention, and webcam-based monitoring fails to provide accurate insights about learners’ mental focus as it is deceived by mere screen fixation without cognitive engagement. Existing BCI-based approaches lack real-time validation and evaluation procedures. To address these limitations, a Brain-Computer Interface (BCI) system is developed using a non-invasive Electroencephalogram (EEG) headband, FocusCalm, to record brainwave activity under attentive and non-attentive states. 20 minutes of data were collected from each of 20 participants watching a pre-recorded educational video. The data validation employed a novel intra-video questionnaire assessment. Subsequently, collected signals were segmented (sliding window), filtered (Butterworth bandpass), and cleaned (removal of high-amplitude and EOG artifacts such as eye blinks). Time, frequency, wavelet, and statistical features were extracted, followed by recursive feature elimination (RFE) with support vector machines (SVMs) to classify attention and non-attention states. The leave-one-subject-out (LOSO) cross-validation accuracy was found to be 88.77%. The system provides feedback alerts upon detection of a non-attention state and maintains focus profile logs. A pilot study was conducted to evaluate the effectiveness of real-time feedback. Five participants underwent a 10-minute session comprising a 5-minute baseline phase devoid of feedback, succeeded by a 5-minute feedback phase, during which alerts were activated if participants exhibited inattention for approximately 8 consecutive seconds. A paired t-test (t = 5.73, p = 0.007) indicated a statistically significant improvement in concentration during the feedback phase.
@@ I Introduction
Concentration monitoring has become essential to ensure learning effectiveness in online classes, whose adoption has been steadily increasing across different educational curricula [1]. Prior research shows that learners are more prone to mind-wandering when they are watching lecture videos (being regarded as online learning scenario) than during live lectures (traditional offline classes) [2]. Thus, to track the engagement of students in educational and learning settings, conventional methods of developing questionnaires [3] or camera-based monitoring [4, 5] are used. Although questionnaires are useful for the objective evaluation of students, they only serve as a manual tool that requires human intervention and input. Conversely, webcam-based monitoring systems are limited to analyzing external behavioral cues such as facial expressions, gaze, or posture, and therefore lack the ability to provide direct neurophysiological insights into learners’ cognitive attention levels [6], alongside inducing privacy issues [7]. Additionally, vision-based models fail to preserve robustness in low-light conditions. For this reason, Brain-Computer Interface (BCI) has emerged as a powerful tool to perform cognitive load evaluation of students while watching lecture videos.
@@
BCI is a form of neurotechnology that is leveraged to establish a direct connection between humans and machines. Specifically, electroencephalography (EEG) signals are increasingly being utilized for various use cases, including bionic intelligence[8], vehicle control [9, 10, 11], and other assistive technologies [12, 13, 14, 15]. BCI has gained considerable recognition in the classification of attentive and non-attentive states [16, 17, 18, 19], including distraction [20] and drowsiness[21]. Despite the fact that this article is particularly oriented to enhance concentration levels, terms like focus, attention, concentration, and engagement, which are different constructs overlapping each other, can be used interchangeably as the goal is to improve online learning for students using BCI. It has been seen from studies that the prefrontal cortex plays a crucial role in evaluating sustained attention [22, 23, 24]. In the frontal region, the features contributing to concentration level detection are predominantly the changes in delta, theta, alpha, and beta band power. [25, 22, 23].
@@
Current state-of-the-art works use EEG devices for laboratory experimentation that are not fully consumer-grade, which makes them not suitable for real-time implementation. Some of them are Emotiv Insight, Emotiv EPOC+, and OpenBCI Ultracortex. These devices lack dry electrodes, requires accurate placement across multiple scalp regions, not cost-effective and present setup challenges for consumers [26, 27]. Among consumer-grade devices, cost-effective options such as Muse, NeuroSky MindWave, and FocusCalm were studied. However, existing works provide limited details about their data collection strategies in online learning scenarios [28, 29, 30, 31], leaving a gap in collecting ground truth (referring to correct labeling of neural data containing actual concentration and actual non-concentration states). Other consumer-grade devices from companies such as Cognionics and Wearable Sensing have been promoted for use in sports scenarios, but their high cost and large number of electrodes limit their suitability for consumer-friendly applications in this field. Vital gaps are seen in real-time implementation of BCI research articulated to track the concentration of students in learning environments. Huang et al. [32] focused on real-time cognitive monitoring to improve concentration time with neuro-feedback mechanism. However, the data collection strategy included mental arithmetic tasks, which is not an identical simulation of a learning environment, and the data validation was done using post regulation assessments such as the Sustained Attention to Response Task (SART) [33], which fails to validate the participants’ engagement during the mental arithmetic tasks. Furthermore, the working algorithm of the single-channel EEG headband, along with the feature extraction and model training process, were not discussed, as their work focused more on evaluating the effects of neurofeedback rather than on developing the classification methodology. Conrad and Newman developed a strategy to use EEG signals for evaluating mind-wandering in online classes [34]. However, the use of only pre- and post-assessment questionnaires is insufficient to validate learners’ attention levels during the video. The work done by Rehman et al. [35] relied on publicly available datasets, which poses the same limitations as the earlier research gaps. The studies summarized in Table I are selected to represent existing approaches for monitoring learner attention in online learning environments. Inclusion criteria are works that addressed learner engagement or attention in online or remote learning settings, and use of other related modalities (e.g., webcam-based, neurofeedback) whose limitations are being addressed in our proposed approach. Exclusion criteria included works focusing purely on offline classroom learning, cognitive tasks unrelated to learning (e.g., driving simulators), or studies that did not explicitly address attention/engagement measurement.
@@
In this study, a eeg based attention classification framework is proposed to address the research gaps, with a focus on real-time implementation in online learning scenarios. A novel data acquisition and validation design through an intra-video questionnaire assessment was done to label the blocks as ground truths. The experimentation was done on classifying two cognitive states of students, namely attention and non-attention. Twenty participants were recruited to participate in data collection experiments containing two sessions, designated for the two classes of data and the device used is a FocusCalm headband. After that, several preprocessing techniques were employed to get clean data, followed by feature extraction, feature selection and classification. Following this, leave-one-subject-out validation was used in order to evaluate the classifier model when faced with unseen data. Finally, a simple app GUI was developed to have real-time insights of concentration levels along with a feedback mechanism to alert the learner during non-attention states. Thus, this article has contributions in the following areas:
  • A data collection paradigm is introduced that uses authentic lecture videos from educational technology platforms as EEG stimuli, offering a closer approximation to real-world online learning environments than traditional controlled tasks.
  • A novel intra-video questionnaire assessment is proposed for data validation, which, unlike traditional post-task questionnaires or external measures (e.g., SART), collects learner responses during the lecture itself. This approach minimizes recall bias, provides time-specific ground-truth labels, and enhances the reliability of the EEG dataset.
  • A real-time system is developed with feedback alerts to ensure the attention spans of learners improve over time. Preliminary evidence of its effectiveness in enhancing concentration is validated via a pilot study.
@@ II Data Collection
The section gives a detailed explanation of the inclusion criteria of participants, the device used for data collection, the experimental design, and setup.
@@ II-A Participant Selection
A total of 20 students with a mean age of $22.8\pm 3.8$ years were recruited for the data collection experiment. A detailed explanation of the experimental setup, collection paradigm, and the study was given to the participants before taking their informed consents. Participants retained the right to leave the data collection process at their discretion at any moment. This data collection protocol was approved by the Institutional Research Ethics Board (IREB), United International University (UIU). Among the participants, 15 are graduate students of Advanced Intelligent Multidisciplinary System Lab (AIMS Lab) under Institute of Research, Innovation, Incubation & Commercialization (IRIIC), UIU and rest are the students of UIU. They were selected to pass the condition of having cognitive abilities necessary to comprehend a standard high school or equivalent curriculum-based educational material. This was ensured due to the video stimuli being exposed to the participants at the time of data collection, which contained online lectures on different topics prepared by national educational technology platforms.
@@ II-B Data Acquisition Device
FocusCalm headband [36] was used as the data acquisition device in this work for its consumer-grade accessibility, dry electrode system, frontal EEG coverage, low setup complexity and a balanced trade-off between signal quality and user comfort. It is a neuro-feedback EEG device designed for meditation purposes. There are three dry electrodes placed close to each other and aligns with the Fp1, Fp2, and Fpz positions of the international 10-10 electrode placement system as shown in Fig. 1. Fpz is used as the primary electrode, while Fp1 and Fp2 serve as the reference and ground electrodes [37]. The device wirelessly sends single channel data at 250 Hz, along with eight distinct values, of which six indicate the delta, theta, alpha, low beta, high beta, and gamma frequency band values, and the other two show attention and meditation scores given by their developed model. The developers of FocusCalm at BrainCoTech have made a software development kit (SDK) named Crimson SDK, which is accessible for taking raw EEG data along with the eight other values in real-time.
@@ II-C Data Collection Experiment
The data collection experiment was designed using the PsychoPy software [38]. The design included a starting instructional video for both classes (attentive and non-attentive) that gives a brief view of the experiment done.
The participants were asked to watch a 10-minute video lecture that was split into 30-second blocks. Subsequently, a question on that exact block was shown on the screen with multiple choices, which the participants had to answer. The duration given for answering the question was set to 15 seconds, as an average human takes around 4 to 5 seconds for reading a sentence [39]. This gives the participants 10 seconds to answer which is considered ample, as humans take around 2 seconds to recognize answers [40]. The questions were set to ensure that the participant has followed the lecturer’s quotes and drawings on the board. The questions require no external knowledge to be answered in order to ensure an objective evaluation of a participant’s engagement. After the instructional video, a short sample video was played followed by a question-answer block to familiarize the participant with the experiment. A conditional screen was displayed with a fixation cross where the participant has to press a key to start the video lecture. In this way, each participant had to watch 20 video blocks of a lecture followed by questions inserted between them. The experimental setup captures timestamps, which mark the start and end of each video block together with participants’ responses to the questions. These records were subsequently used for data validation.
For the non-attentive data collection session, similar video sequences were presented, but rendered inaudible and blurred to prevent participants from perceiving the content. The participants were instructed to stare at the screen for 5 minutes followed by 5 minutes with eyes closed. In the eyes-open condition, participants viewed the hazy, non-informative video, which was designed as a monotonous, low-load task. Such tasks, which provide minimal external stimulation, are known to induce states of boredom or an ”unengaged mind”, creating a strong tendency for attention to shift from the external environment to internal thought [41, 42]. This method is based on extensive findings that mind-wandering is most prevalent when cognitive demands are low [43], thereby simulating a common occurrence in real-world scenarios where students face difficulties in maintaining attention during online lectures. The eyes-closed condition was designed to simulate drowsiness or dozing off during online lectures. A brief overview is illustrated in Fig. 2a.
@@ II-D Data Validation
After successful data acquisition, the EEG data were stored with proper timestamps through some modifications of the Crimson SDK, and the responses to the questions were stored with synchronized timestamps. The correct response from a participant validates the EEG data corresponding to the video block that follows the question. All validated blocks were stored as ground truths for attentive data. For non-attentive data, the questionnaire assessment was not performed and thus required no further validation. The procedure is shown in Fig. 2b.
@@ III Methods
The validated data were processed through several steps, namely data pre-processing, feature extraction, and feature selection before being used as training data for classification.
@@ III-A Data Preprocessing and Preparation
Data were pre-processed through steps that include segmentation, filtering, noise removal, and smoothing.
@@ III-A1 Segmentation
EEG data of each participant of a total length of $T$ samples were partitioned to get data segments of optimal length $L=1750$ samples (details in section III.D) . Although the average sampling frequency of 250 Hz was retained, samples were used instead of time intervals to avoid errors caused by minute temporal delays in packet transmission, which complicate timestamp-based calculations. Through validation, a total of $K$ blocks were selected from the initially acquired $n$ blocks for each participant. This number corresponds to the average number of questions (around 20) set for a video lecture per session. A sliding window was used during segmentation to reduce the edge effects and improve statistical reliability [44]. The sliding window is also referred to in this article as the overlap ratio $r=0.7$ (details in section III.D). The total number of segments found for each selected block is:
@@
The total number of samples is $x=n\times L$. During segmentation, two columns were added for labeling and assigning subject-wise user_ids. These were used later in classification.
To ensure subject-wise label balance, samples for both classes were extracted separately for each subject. Let $\mathcal{D}_{u}^{0}$ and $\mathcal{D}_{u}^{1}$ denote the sets of samples with labels 0 and 1 respectively (0 meaning attentive and 1 meaning non-attentive), for a given user $u$. An equal number of samples was retained by trimming the larger class to match the size of the smaller one. Number of samples to be retained, $n_{u}=\min(|\mathcal{D}_{u}^{0}|,|\mathcal{D}_{u}^{1}|)$, and the last $n_{u}$ samples from each label set were selected, which gave $\mathcal{D}_{u,\text{tail}}^{0}(n_{u})$ and $\mathcal{D}_{u,\text{tail}}^{1}(n_{u})$. The balanced dataset for user $u$ was then
@@
Given that the dataset retained a substantial number of segments, class imbalance was addressed by trimming the larger class rather than oversampling the smaller one.